# ChronoSpike: An Adaptive Spiking Graph Neural Network for Dynamic Graphs

### v1

@@ Abstract
Dynamic graph representation learning requires capturing both structural relationships and temporal evolution, yet existing approaches face a fundamental trade-off: attention-based methods achieve expressiveness at $O(T^{2})$ complexity, while recurrent architectures suffer from gradient pathologies and dense state storage. Spiking neural networks offer event-driven efficiency but remain limited by sequential propagation, binary information loss, and local aggregation that misses global context. We propose ChronoSpike, an adaptive spiking graph neural network that integrates learnable LIF neurons with per-channel membrane dynamics, multi-head attentive spatial aggregation on continuous features, and a lightweight Transformer temporal encoder, enabling both fine-grained local modeling and long-range dependency capture with linear memory complexity $O(T\cdot d)$. On three large-scale benchmarks, ChronoSpike outperforms twelve state-of-the-art baselines by 2.0% Macro-F1 and 2.4% Micro-F1 while achieving $3\text{--}10\times$ faster training than recurrent methods with a constant 105K-parameter budget independent of graph size. We provide theoretical guarantees for membrane potential boundedness, gradient flow stability under contraction factor $\rho<1$, and BIBO stability; interpretability analyses reveal heterogeneous temporal receptive fields and a learned primacy effect with 83–88% sparsity.
@@ 5.1.3 Implementation Details
Table 3: Ablation study of ChronoSpike for node classification on DBLP, Tmall, and Patent datasets. Tr. Ratio denotes the training ratio. For ChronoSpike, we report mean ± standard deviation over five runs. Ablation results are reported as single runs, following common practice. The symbol “-” denotes under-compilation.
@@ 5.2.2 Ablation Study
We conduct ablation studies on DBLP, Tmall, and Patent under different training ratios to assess the contribution of core components of ChronoSpike (Table 3). Removing the temporal Transformer causes the largest drop, with Micro- and Macro-F1 decreasing by 3-5%, confirming its role in modeling long-range temporal dependencies beyond local spiking dynamics. Replacing adaptive LIF neurons with fixed ones degrades performance, showing the importance of learnable membrane dynamics for capturing complex temporal patterns. Substituting attentive aggregation with mean pooling leads to a moderate but consistent drop, especially on dense Tmall, indicating the benefit of adaptive neighbor weighting. The contrastive loss provides consistent regularization gains across settings. ChronoSpike outperforms both the ReLU-based continuous variant (w/o SNN) and the LSTM baseline, supporting the combination of sparse spiking dynamics with a global Transformer readout. The static baseline performs much worse, particularly on DBLP (as low as 4.56%), showing that dynamic graph learning requires explicit temporal modeling. Full results are reported in Appendix F.1. We further analyze how ChronoSpike captures temporal dependencies via interpretability studies in Appendix F.3.
@@ F.1 Details of Ablation Study
We analyze ChronoSpike on the DBLP and Tmall datasets with different training ratios to validate the contribution of each core component to its performance. We systematically removed or replaced the key modules, as shown in Table 3. The results show that removing the temporal transformer results in the largest drop in performance, with both Micro and Macro F1 scores decreasing by about 3-5%. This component captures global long-range dependencies that local spiking dynamics alone cannot address. Similarly, the removal of adaptive LIF neurons consistently led to worse results. This confirms that learnable membrane time constants and dynamics are crucial to effectively model complex temporal patterns. Further ablations show that the model is not immune to other design choices. Removing the attentive aggregator and using mean pooling caused a moderate drop, particularly on the dense Tmall graph. This highlights the value of adaptive neighbor weighting in this study. The contrastive loss provides a consistent regularization benefit. Crucially, ChronoSpike consistently outperformed the ReLU-based continuous variant (w/o SNN) and the LSTM baseline. This indicates that the combination of sparse spiking dynamics, which are effective at filtering noise, and a global transformer readout performs better than standard continuous activations or recurrent propagation. Finally, the complete failure of the static baseline, particularly on DBLP, where the performance drops to as low as 4.56%, confirms that effective dynamic graph learning fundamentally requires temporal modeling.

### v2

@@ Abstract
Dynamic graph representation learning requires capturing both structural relations and temporal evolution, yet existing approaches face a core trade-off: attention-based methods offer expressiveness at $O(T^{2})$ complexity, while recurrent architectures suffer from gradient pathologies and dense state storage. Spiking neural networks provide event-driven efficiency but are constrained by sequential propagation, binary information loss, and local aggregation that lacks global context. We propose ChronoSpike, an adaptive spiking graph neural network that integrates learnable LIF neurons with per-channel membrane dynamics, multi-head spatially-attentive aggregation over continuous features, and a lightweight Transformer temporal encoder. This design enables fine-grained local modeling and long-range dependency capture with linear memory complexity $O(T\cdot d)$. ChronoSpike outperforms twelve state-of-the-art baselines on three large benchmarks by 2.0% Macro-F1 and 2.4% Micro-F1 on average while achieving $3\text{--}10\times$ faster training than recurrent methods with a constant 105K-parameter budget independent of graph size. We provide theoretical guarantees for membrane potential boundedness, gradient flow stability under contraction factor $\rho\!<\!1$, and BIBO stability; interpretability analyses reveal heterogeneous temporal receptive fields and a learned primacy effect with 83–88% sparsity.
@@ 5.1.3 Implementation Details
Table 3: Ablation study of ChronoSpike for node classification on DBLP, Tmall, and Patent datasets. Tr. Ratio denotes the training ratio. For ChronoSpike, we report mean ± standard deviation over five runs. Ablation results are reported as single runs, following common practice.
@@ 5.2.2 Ablation Study
We conduct ablation studies on DBLP, Tmall, and Patent under different training ratios to assess the contribution of core components of ChronoSpike (Table 3). Removing the temporal Transformer causes the largest drop on DBLP and Tmall (3–5% Micro- and Macro-F1) and a consistent drop on Patent ($\approx$2.7–3.1%), confirming its role in capturing long-range temporal dependencies. Replacing adaptive LIF neurons with fixed ones degrades performance across all benchmarks, showing the importance of learnable membrane dynamics. Substituting attentive aggregation with mean pooling results in moderate drops, most notably on dense Tmall and Patent, where it yields among the lowest component-level ablation scores. The contrastive loss provides consistent regularization gains across all settings. The identity of the strongest ablation variant shifts across datasets, the ReLU variant and contrastive loss ablation on DBLP and Tmall, the LSTM temporal variant on Patent, yet ChronoSpike outperforms all ablations uniformly, with margins of 2.0–2.5 percentage points over the best ablation on Patent, demonstrating that sparse spiking dynamics and Transformer-based temporal aggregation provide complementary benefits. The static baseline degrades substantially everywhere, most dramatically on DBLP (as low as 4.56%) and by 5–6% on Patent, confirming that temporal modeling is indispensable regardless of dataset scale. Full results are reported in Appendix F.1. We further analyze how ChronoSpike captures temporal dependencies via interpretability studies in Appendix F.3.
@@ F.1 Details of Ablation Study
We analyze ChronoSpike on the DBLP, Tmall, and Patent datasets with different training ratios to validate the contribution of each core component to its performance. We systematically removed or replaced the key modules, as shown in Table 3. Removing the temporal Transformer results in the largest performance drop on DBLP and Tmall, with both Micro and Macro F1 scores decreasing by about 3–5%, confirming that this component captures global long-range dependencies that local spiking dynamics alone cannot address. On Patent, the drop is smaller but still consistent ($\approx$2.7–3.1%), reflecting that its shorter temporal horizon (T=25) reduces the relative advantage of global temporal attention; nonetheless, the Transformer remains beneficial across all settings. Similarly, the removal of adaptive LIF neurons consistently led to worse results on all three datasets, confirming that learnable membrane time constants are crucial to effectively model complex temporal patterns. Removing the attentive aggregator and using mean pooling caused a moderate but consistent drop across datasets, with the effect most pronounced on the dense Tmall graph. On Patent, the attention ablation yields among the lowest component-level scores at most training ratios (e.g., 0.8240 Mi-F1 at 80%), though at 60% the ReLU variant scores slightly lower, suggesting that both spatial attention and spike-based computation contribute meaningfully on large-scale graphs. The contrastive loss provides a consistent regularization benefit across all three datasets and training ratios.
The identity of the strongest ablation variant varies across datasets and settings. On DBLP, the ReLU variant (w/o SNN) and the adaptive LIF ablation alternate as the top-performing ablations; on Tmall, the contrastive loss ablation and the ReLU variant are frequently among the strongest. On Patent, the LSTM temporal variant performs best across all training ratios (e.g., 0.8403 Ma-F1 and 0.8391 Mi-F1 at 80%), suggesting that Patent’s large spatial scale and shorter temporal depth make recurrent temporal modeling a competitive alternative. Despite these variations, ChronoSpike consistently outperforms all ablation variants across every dataset and training ratio, with margins of approximately 2.0–2.5 percentage points over the best ablation on Patent, indicating that the combination of sparse spiking dynamics and Transformer-based temporal readout provides complementary benefits beyond what any single component achieves individually. Finally, the static baseline confirms that effective dynamic graph learning fundamentally requires temporal modeling. The failure is most dramatic on DBLP, where performance drops to as low as 4.56%, while on Patent, the degradation is less severe (Ma-F1 $\approx$0.805), likely because Patent’s rich node features carry substantial class-discriminative information even without temporal context. Nonetheless, the consistent gap of roughly 5–6% on Patent and the near-complete collapse on certain DBLP splits confirm that temporal modeling is indispensable.